OpenStudy (anonymous):

square root property (x+2)^2=-7

OpenStudy (danjs):

\[\sqrt{-1}=i\]

OpenStudy (danjs):

\[\sqrt{-7}=\sqrt{-1*7}=\sqrt{-1}\sqrt{7}=\sqrt{7}*i\]

OpenStudy (loser66):